OpenStudy (genny7):

Choose the two equations you would use to solve the absolute value equation below. Then solve the two equations. |x + 5| = 55 A. x + 5 = 55 and x + 5 = –55; {50, –60} B. x + 5 = 55 and x + 5 = –55; {–50, –60} C. x – 5 = 55 and x + 5 = –55; {60, –60} D. x – 5 = 55 and x + 5 = –55; {–60, –60}

OpenStudy (vishweshshrimali5):

What do you think the answer must be ?

OpenStudy (genny7):

i think its C

OpenStudy (vishweshshrimali5):

No its not C. Remember this: \[\large{|x+a| = b}\] \[\large{\implies x + a = \pm b}\] \[\large{\implies x + a = b~~\text{or}~~x+a = -b}\]

OpenStudy (vishweshshrimali5):

Now what do you think the answer should be ?

OpenStudy (genny7):

A?

OpenStudy (vishweshshrimali5):

Fantastic !!

OpenStudy (vishweshshrimali5):

You are correct :)

OpenStudy (genny7):

thanks sooo much! :)

OpenStudy (vishweshshrimali5):

Your welcome ;)

OpenStudy (genny7):

could you help me on another question?

OpenStudy (vishweshshrimali5):

Yeah sure but please post it as a new question and tag me there :)

OpenStudy (genny7):

ok